### 05 Mar 2026 v1

# AutothinkRAG: Complexity-Aware Control of Retrieval-Augmented Reasoning for Image-Text Interaction

Abstract
Information-intensive Document Question Answering (DocQA) is often constrained by long contexts and information overload, which hinders Vision-Language Models (VLMs) from performing precise direct reasoning. Although multimodal GraphRAG has achieved preliminary breakthroughs, existing frameworks still face dual challenges: (1) the necessity of large-scale models for handling queries of diverse complexities and (2) the inherent reasoning bottlenecks of end-to-end VLMs. To address these issues, we propose AutoThinkRAG, a framework that enhances the understanding of complex documents by synergizing the capabilities of multiple models. Specifically, we introduce a Query Complexity Router to allocate reasoning paths based on the analysis of query difficulty. Furthermore, to overcome the reasoning boundaries of VLM, we propose a functional decoupling architecture: a small-scale VLM serves as a high-fidelity visual interpreter to transform query-relevant visual cues into textual representations, which are subsequently processed by an LLM for logical deduction and synthesis. Extensive experiments on DocBench and MMLongBench demonstrate that AutoThinkRAG significantly reduces inference costs while achieving new state-of-the-art performance. Further ablation studies verifies the effectiveness of our proposed method.
AutothinkRAG: Complexity-Aware Control of Retrieval-Augmented Reasoning for Image-Text Interaction
Jiashu Yang1,3, Chi Zhang1, Abudukelimu Wuerkaixi2,3, Xuxin Cheng3, Cao Liu3, Ke Zeng3, Xu Jia1, Xunliang Cai3 1Dalian University of Technology 2Tsinghua University 3Meituan LongCat Interaction Team
1 Introduction
Retrieval-Augmented Generation (RAG) serves as a core technology for processing heterogeneous multimodal documents, including financial PDFs and technical diagrams. Its objective is to retrieve multimodal information relevant to a given query from documents as contextual content for question answering, thereby enhancing the accuracy and richness of generated results. As an evolving direction in this field, Multimodal RAG (MM-RAG) needs to address two core issues simultaneously: multimodal information processing and high-precision reasoning.
Knowledge graphs can effectively represent the relationships and attributes of information, while leveraging dense vector representations to identify semantically relevant content lacking direct structural connections. In recent years, researchers have conducted extensive explorations on multimodal graph RAG and complex question answering. As the state-of-the-art (SOTA) approach, RAGAnything (Guo et al., 2025) has validated the application of multimodal graphs in document question answering. HyperGraphRAG (Luo et al., 2025) effectively captures high-order dependencies in complex knowledge bases through hyperedge modeling, improving the accuracy of complex multi-hop reasoning. MMGraphRAG (Wan and Yu, 2025) provides a new technical paradigm for multimodal RAG in graph RAG scenarios.
Despite the progress achieved by existing methods, current SOTA frameworks still face two critical issues that severely impact performance:(1) Retrieval Rigidity: Existing systems adopt a static retrieval strategy that fails to make reasonable judgments and planning based on query complexity. Accurate understanding of queries with arbitrary complexity requires large-scale models, and this high-computation-cost approach leads to the inefficient allocation of computing resources.(2) Reasoning Deficit: Multimodal information question answering relies on monolithic Vision-Language Models (VLMs) for end-to-end generation. However, recent related studies have shown that VLMs exhibit significantly lower reasoning performance than Large Language Models (LLMs), which consequently results in the problem of "correct visual recognition but incorrect answer generation"(Wan and Yu, 2025).
In this paper, we propose the AutoThink-RAG framework, which addresses the aforementioned issues through modular cognitive orchestration. To tackle retrieval rigidity, we design the AutoThink lightweight cognitive router, which dynamically selects retrieval paths based on query parsing and complexity evaluation, enabling on-demand allocation of computing resources. To resolve reasoning deficit, we propose the Decomposition of Perception and Reasoning (DPR) architecture, which limits the role of VLMs to that of "visual translators" and delegates logical integration tasks to LLMs, achieving more accurate reasoning without loss of relevant information.
Our contributions are three-fold:
  • We present AutoThink-RAG, a scalable architecture that integrates MinerU-based parsing with a hybrid Graph-Vector storage, establishing a new pareto-optimal frontier between efficiency and accuracy;
  • We design the AutoThink Router to handle queries of unknown complexity; by leveraging a lightweight Small Language Model (SLM) for in-depth complexity analysis and task decomposition, this module enables adaptive execution path selection, effectively addressing the challenge of Retrieval Rigidity.
  • We introduce a decoupled paradigm for multimodal problem solving, which explicitly separates information transformation from the reasoning phase. This approach addresses the limitations of traditional methods that rely on end-to-end direct inference using VLMs.
  • Experiments on two document-understanding benchmark display that AutoThinkRAG achieves the new state-of-the-art performance without relying on large-scale models.
2 Related Work
2.1 Graph-based Retrieval-Augmented Generation
Traditional retrieval-augmented generation (RAG) relies on chunking and vector retrieval, yet often falls short in dispersed evidence localization and multi-hop reasoning over long and multi-document contexts. GraphRAG (Edge et al., 2024a) addresses this by modeling relations among retrieval units as graphs to enable structured retrieval and reasoning. Subsequent work improves (i) efficiency and evidence coverage via lightweight/ephemeral graphs and multi-stage retrieval (LightRAG (Guo et al., 2024), REANO (Fang et al., 2024a), HopRAG (Liu et al., 2025)) and (ii) multi-hop evidence organization by extracting candidate subgraphs/paths and forming structured evidence chains (GNN-RAG (Mavromatis and Karypis, 2024), TRACE the Evidence (Fang et al., 2024b), PathRAG (Chen et al., 2025a)). Hierarchical designs further support multi-granularity reasoning (ArchRAG (Wang et al., 2025a), From Local to Global (Edge et al., 2024b)), while robustness is strengthened through subgraph pruning, subgraph-level retrieval, or external graph knowledge (G-Retriever (He et al., 2024), RAGRAPH (Jiang et al., 2024)). Beyond binary relations, hypergraph and higher-arity modeling improves expressiveness and reduces information loss in complex reasoning (HyperGraphRAG (Luo et al., 2025), Structure Is All You Need (Lee and Whang, )).
2.2 Multimodal Retrieval-Augmented Generation
Multimodal RAG augments text-based RAG with images and videos as external knowledge sources, mitigating limitations of parametric memory (RA-TTA (Lee et al., 2025), MMed-RAG (Xia et al., 2024)); however, many systems still depend on textification or single-modality alignment, hindering the preservation of fine-grained cross-modal structures. For visually rich documents, full-page visual retrieval units avoid OCR/linearization loss (VisDOM (Suri et al., 2025), VisRAG (Yu et al., 2024), ColPali (Faysse et al., 2024)). In image captioning and VQA, robustness issues (e.g., retrieval instability and frequency bias) are commonly mitigated via diversified retrieval and late-interaction mechanisms. For long videos, segment-level or query-aware retrieval reduces computational cost (VideoRAG (Ren et al., 2025)). Retrieval also benefits vision–language contrastive learning and generation for improved generalization and alignment (retrieval-enhanced VLMs (Iscen et al., 2023), GarmentAligner (Zhang et al., 2024)); in domains such as healthcare and content moderation, it introduces high-confidence or personalized priors at the expense of added system complexity (Sun et al., 2025).
2.3 Document Understanding with Retrieval-Augmented Generation
Long- and multi-document QA is constrained by context length and evidence localization difficulty. Recent methods improve retrieval granularity and context packaging through logical block decomposition and iterative retrieval (Document Haystacks (Chen et al., 2025b)), document restructuring and adaptive merging (Virtual Documents (Mass et al., 2024), Document Segmentation Matters (Wang et al., 2025c)), and structured indexing or hierarchical summarization (PdfTriage (Saad-Falcon et al., 2024), MemSum-DQA (Gu et al., 2023), BookRAG (Wang et al., 2025b)). Classical paradigms establish coarse-to-fine localization (Dynamic Coattention Networks (Xiong et al., 2016), Coarse-to-Fine QA (Choi et al., 2017)), and later work models global structure and discourse hierarchies (Compressive Graph Selector Networks (Nie et al., 2022), Beyond Chunking (Chen et al., 2026)); for cross-document reasoning, link-aware pretraining further strengthens inter-document semantic modeling (Yasunaga et al., 2022). Retrieval and generation are increasingly coupled via iterative/corrective schemes (Generate-then-Ground (Shi et al., 2024), REAR (Wang et al., 2024), Active RAG (Jiang et al., 2023), Iterative Retrieval–Generation Synergy (Shao et al., 2023), IAG (Zhang et al., 2023)). For visually rich documents, multimodal document RAG incorporates layout, entities, and knowledge structures for multi-page understanding (REVEAL (Hu et al., 2023), VDocRAG (Tanaka et al., 2025), VisDoM (Suri et al., 2025), MMGraphRAG (Wan and Yu, 2025)); agent-based methods further improve robustness via multi-agent collaboration and iterative feedback (DocAgent (Yang et al., 2025), DocLens (Zhu et al., 2025), Doc-React (Wu et al., 2025)).
3 Methodology
In this section, we first present the overall workflow of AutoThinkRAG. Subsequently, we specify the strategies for information extraction, storage, and transmission. Finally, we detail the implementation of the Query Complexity Router and the Functional Decoupling Architecture.
3.1 Overall Workflow
AutoThinkRAG deconstructs the multimodal Document Question Answering (DocQA) task into intent-aware routing and decoupled reasoning. The end-to-end inference process is formally defined as:
$A=\Psi\left(Q,\mathcal{D}\right)=\text{LLM}\left(T_{v},\mathcal{R},I_{p}\right)$ (1)
where $Q$ is the input query, $\mathcal{D}$ is the document collection, $\mathcal{R}$ denotes retrieved relevant context, $T_{v}$ is the textualized visual evidence, $I_{p}$ is the reasoning path instruction, and $A$ is the final synthesized answer. The workflow proceeds through three core stages:
  1. 1.
    Knowledge Base Construction: The system parses $\mathcal{D}$ into content blocks $b$ with metadata $\mathcal{M}_{b}=\{T,C,B,P,S\}$, encompassing type, content, spatial coordinates, page, and storage path. These are integrated into a hybrid Graph Knowledge Base (GKB) and vector store to support spatial-semantic retrieval (detailed in Sec. 3.2).
  2. 2.
    Query Complexity Routing (QCR): For each query $Q$, the QCR performs a pre-execution analysis to output a complexity label $c(Q)\in\{\text{Simple, Moderate, Complex}\}$ and routing instructions $I_{p}$. This stage determines sub-query decomposition and optimizes retrieval paths for the recall module $\mathcal{R}$ (detailed in Sec. 3.3.1).
  3. 3.
    Decomposition of Perception and Reasoning (DPR): Guided by $I_{p}$, the system fetches context $\mathcal{R}$ and invokes raw visual assets via path $S$. A training-free small-scale VLM serves as a visual interpreter to generate structured textual descriptions $T_{v}$. Finally, a high-performance LLM integrates $\{T_{v},\mathcal{R}\}$ to synthesize the final answer $A$ through rigorous deduction following $I_{p}$ (detailed in Sec. 3.3.2).
By functionally decoupling perception from reasoning, AutoThinkRAG bridges the logical bottlenecks in standalone models, ensuring robust comprehension across extensive multimodal contexts.
3.2 Information Extraction, Storage, and Transmission
To support complex multimodal reasoning, we design a high-fidelity information processing pipeline. This pipeline is engineered to transform heterogeneous documents (e.g., PDF, PPT) into a structured and queryable knowledge system while preserving the original spatial layout and semantic integrity.
3.2.1 High-Fidelity Parsing and Metadata Extraction
We employ a robust parsing engine (e.g., Mineru) to perform deep analysis on the input document $D$. The parsing output comprises a structured JSON file containing decomposed content blocks $\mathcal{B}$ and an asset repository for multimodal objects (images, tables, equations). For each extracted block $b\in\mathcal{B}$, we maintain a metadata tuple $\mathcal{M}_{b}=\{T,C,B,P,S\}$, where:
  • $T$ denotes the Type (e.g., text, image, or table);
  • $C$ represents the raw Content or the caption associated with the block;
  • $B$ is the Bounding Box (Bbox) coordinates for spatial localization;
  • $P$ indicates the Page Number;
  • $S$ is the Storage Path, serving as the unique index for asset transmission.
3.2.2 Differentiated Information Processing
The system implements bifurcated semantic extraction logic based on content type:
  • Textual Stream: Raw text is serialized and partitioned into chunks using a sliding window of 1,200 tokens with a 100-token overlap. Subsequently, an Information Extraction (IE) module identifies entity and relation triplets from each chunk.
  • Multimodal Stream: For non-textual assets, the system generates a detailed semantic Description. The original Caption is concatenated with this description to form an augmented representation, followed by IE to align visual concepts with the global knowledge schema.
3.2.3 Hybrid Storage and Transmission Protocol
Processed information is integrated into a dual-layered storage architecture:
  1. 1.
    Graph Knowledge Base (GKB): We employ a hard-matching algorithm for entity disambiguation and merging to construct a global relational graph.
  2. 2.
    Vector Store: All entities, relations, and text chunks are mapped into a dense embedding space, with $S,B,$ and $P$ indexed as key attributes.
Information Transmission Mechanism. The framework achieves efficient cross-module transmission via a metadata-driven protocol. During retrieval, the storage path $S$ enables the direct transmission of raw multimodal data to the VLM. Simultaneously, leveraging $B$ and $P$, the system fetches adjacent context from the original document, effectively bridging the “Information Gap” between isolated fragments and their original document context.
Doc Types Domains All
Aca. Fin. Gov. Law. News Txt. Mm. Una.
LLM 41.3 16.3 50.7 49.7 77.3 53.9 20.1 75.8 44.8
MLLM 19.8 16.3 28.4 31.4 46.5 35.7 15.9 39.5 27.7
$M$ Locations Formats Overall
Sin. Mul. Una. C.T. Txt. Lay. Fig. All
LLM 22.5 20.0 53.2 14.7 33.3 23.5 16.1 27.8
MLLM 13.3 7.9 13.9 8.4 10.7 11.8 11.4 11.6
Table 1: Comparative experimental results of VLM and LLM on DocBench and MMLongBench benchmarks.
Figure 1: Overview of our AutoThink–RAG. The framework consists of four primary stages: (1) Information Extraction: Heterogeneous multimodal data are ingested. (2) Data Processing: Through layout analysis, the system partitions data into text chunks and image indices, followed by an iterative “Extract–Judge” loop to construct a Graph Knowledge Base. (3) Data Linker: A hybrid retrieval mechanism integrates traditional Top–K vector search with graph–based merging to capture complex entity relationships. (4) Reasoning & Response: The Query Router decomposes the original query into sub-queries ($Q_{1},Q_{2},Q_{3}$), which are then processed through a decoupled perception reasoning architecture leveraging both “related text” and visual cues to synthesize the final response.
3.3 AutoThink Router and Functional Decoupling Architecture
This section details the design motivation and technical implementation of the Query Complexity Router (QCR) and the Decomposition of Perception and Reasoning (DPR) strategy.
3.3.1 Query Complexity Router (QCR)
The fidelity of RAG systems is fundamentally constrained by the coupling between query intents and corresponding contexts. Given the sequential nature of the pipeline, query comprehension serves as a critical bottleneck. When encountering queries of unknown complexity, conventional end-to-end approaches often necessitate large-scale models to parse implicit meanings or nested issues, incurring high costs. The QCR addresses this by directly assessing query complexity to influence subsequent retrieval and processing paths, thereby indirectly bolstering model accuracy. We employ an SLM to determine whether a query warrants decomposition into sub-questions, routing it to optimal execution paths at minimal computational cost.
Technical Implementation: The router extracts three categories of features from $Q$:
  1. (a)
    Semantic Feature: Captures core intent (e.g., factual retrieval, comparative reasoning, causal analysis);
  2. (b)
    Element Feature: Counts entities $|\mathcal{E}(Q)|$, visual references $|\mathcal{M}(Q)|$, and constraints;
  3. (c)
    Dependency Feature: Identifies cross-chunk/multi-step reasoning needs ($\text{ctx}(Q)$).
Based on these, QCR assigns a complexity label $c(Q)$ and generates routing instructions $I_{p}$.
3.3.2 Functional Decoupling via DPR
Recent studies indicate that VLMs exhibit significantly lower capabilities in logical reasoning compared to LLMs (as shown in Table 1). This discrepancy stems from the VLM architecture’s over-reliance on intermediate projection layers for modality alignment, which struggles to maintain logical consistency in complex derivations. Conversely, VLMs demonstrate superior performance in descriptive tasks post-perception. Consequently, we propose the Decoupled Perception and Reasoning (DPR) strategy, partitioning the task into two specialized sub-processes:
  1. 1.
    Visual Perception (Small-scale VLM): A lightweight VLM (e.g., Qwen2.5-VL-3B) serves as a visual interpreter, generating a structured description $T_{v}$ in a training-free, zero-shot mode:
    $T_{v}=\text{VLM}_{\text{small}}(v,P)$ (2)
  2. 2.
    Logical Reasoning (LLM): Multimodal DocQA is transformed into text-only reasoning. A unified context $C=\{T_{v}\}\cup\mathcal{R}(Q,\mathcal{D})\cup\{\text{Intent}(Q)\}$ is constructed. The LLM then synthesizes the answer $A$ guided by path-specific instructions:
    • Simple: “Extract target row/column value from table descriptions.”
    • Moderate: “Calculate/compare data using table descriptions (show steps).”
    • Complex: “Integrate multi-table/text data for step-by-step analysis.”
    $A=\text{LLM}(C,\text{Instruction}(c(Q)))$ (3)
By decoupling perception from reasoning, the LLM performs rigorous deduction over precise textualized visual evidence, effectively bridging the reasoning gap of standalone multimodal models.
Method Domains Types Overall
Aca. Fin. Gov. Law. News. Txt. Mm. Una.
AutothinkRAG 80.16 75.40 85.71 87.09 92.50 86.43 77.31 81.25 82.13
Nothink 77.34 73.46 80.59 85.29 88.78 84.08 76.12 77.84 79.37
VLM only 77.03 74.00 81.54 84.39 89.26 85.32 75.09 76.42 79.47
baseline 76.34 72.93 79.77 84.07 84.52 84.65 74.47 75.83 78.02
Raganything 74.90 70.68 77.41 83.06 81.67 84.21 76.36 52.80 75.47
Table 2: Accuracy (%) on DocBench Dataset. Domain categories include Academia (Aca.), Finance (Fin.), Government (Gov.), Legal Documents (Law), and News Articles (News). Document types are categorized as Text-only (Txt.), Multimodal (Mm.), and Unanswerable queries (Una.).
Method Domains Overall
Res. Tut. Acad. Guid. Broch. Admin. Fin.
AutothinkRAG 44.86 52.20 52.26 52.34 51.00 62.19 56.14 51.29
Nothink 43.15 48.55 43.21 52.25 46.00 53.08 50.42 47.21
VLM only 41.43 49.27 43.43 52.25 41.00 54.32 48.72 45.55
baseline 40.41 47.82 42.42 50.96 42.00 51.85 46.15 44.86
8B 42.80 44.12 43.21 52.34 46.00 50.61 50.42 46.09
Raganything 41.09 41.30 44.22 45.16 40.00 54.32 52.13 44.36
Table 3: Accuracy (%) on MMLongBench across different domains and overall performance. Domain categories include Research Reports/Introductions (Res.), Tutorials/Workshops (Tut.), Academic Papers (Acad.), Guidebooks (Guid.), Brochures (Broch.), Administration/Industry Files (Admin.), and Financial Reports (Fin.).
4 Experiments
In this section, we empirically validate the effectiveness of AutoThink-RAG. Our goal is to address two primary research questions: (1) Compared to static retrieval, can the AutoThink router effectively reduce hallucinations and handle complex query intents? (2) Does the Decoupled Perception-Reasoning (DPR) architecture demonstrate superior performance in reasoning tasks involving multimodal documents of varying lengths?
4.1 Experimental Setup
Benchmarks.
We evaluate our framework on two comprehensive multimodal benchmarks:
  • DocBench : A diverse dataset covering five domains (Academic, Financial, Government, Law, News) and three query types (Text-only, Multimodal, and Unanswerable). The "Unanswerable" category is particularly critical for assessing model hallucination and refusal capabilities.
  • MMLongBench : A benchmark designed to test long-context understanding across varied document types (e.g., Research Papers, Manuals, Financial Reports), requiring the model to synthesize information across multiple pages and modalities.
4.2 Implementation Details
Figure 2: Comparison of routing strategy distributions across different document lengths. This figure illustrates the selection proportions of graphs and hypergraphs within length intervals under Think and No-Think modes.
In our experimental evaluation of AutoThink-RAG, we implement a multi-model collaborative framework characterized by a decoupled architecture. Specifically, we employ Qwen2.5-VL-3B as the AutoThink Router to facilitate global task orchestration and intent scheduling. For the knowledge construction phase, Qwen3-8B (non-reasoning version) is utilized to perform robust entity and relation extraction. To optimize retrieval precision, we adopt Qwen3-embedding-8B (with 1024-dimensional embeddings) for dense vector representation, followed by bge-reranker-v2-m3 for fine-grained reranking. To realize the Decoupled Perception-Reasoning (DPR) paradigm, we deploy Qwen2.5-VL-7B as the Perception Expert, which distills complex visual information into dense textual descriptions. Subsequently, Qwen3-32B-Instruct serves as the Reasoning Expert to execute high-level logical deductions based on the integrated context. Regarding hyperparameter configurations, all generative models are set with a temperature of $0.7$ and a Top-$p$ value of $0.9$. During the retrieval stage, we set the Top-$K$ for entity recall to $40$ and the number of retrieved text chunks to $20$ to ensure sufficient context coverage.
4.3 Main Results
Performance on DocBench.
Table 2 summarizes the performance on DocBench. AutoThink-RAG achieves a state-of-the-art Overall Accuracy of 82.13%, significantly outperforming the baseline (78.02%).
Crucially, our framework demonstrates a substantial breakthrough in the Unanswerable (Una) category, boosting accuracy from 52.80% to 81.25% (a +28.45% improvement). This result strongly validates the cognitive awareness of our AutoThink router. Unlike static retrieval systems that force-feed irrelevant context to the generator, AutoThink can detect insufficient information and guide the reasoning engine to refuse answering, thereby drastically reducing hallucinations. Additionally, in information-dense domains like News (+10.83%) and Government (+8.30%), our Hybrid Hypergraph Retrieval proves effective in capturing complex entity relationships that simple vector search misses.
Figure 3: The approach of converting VLM outputs to LLM leads to a marked boost in accuracy when processing multi-page long documents.
Performance on MMLongBench.
Table 3 presents the results on long-context tasks. AutoThink-RAG consistently surpasses the baseline across all document categories, achieving an Overall Accuracy of 51.29% (+6.43%).
The unified baseline often struggles with "contextual entrainment" in long documents, where visual noise distracts the VLM’s reasoning. In contrast, our DPR Architecture maintains high performance even in complex categories like Admin (+10.34%) and Finance (+9.99%). By converting visual signals into dense text descriptions and offloading the synthesis to a specialized LLM, our approach effectively bridges the "Reasoning Gap," allowing for robust multi-hop deduction over extensive contexts.
4.4 Baselines
To ensure experimental rigor and fairness, we built a baseline model and employed the identical model on RAGAnything (Guo et al., 2025) for testing and comparison.In the baseline configuration, we bypass query decomposition and employ the VLM for direct end-to-end inference, rather than using our proposed decoupled paradigm. For the RagAnything benchmark, we substitute its internal components with the specific models used in our framework and maintain identical hyperparameter configurations to ensure a controlled comparison.
4.5 Ablation Study
To demonstrate the superiority of the proposed architecture, we conducted ablation studies by systematically removing the routing and decoupling modules, respectively. Furthermore, we performed an in-depth analysis of relevant performance metrics across varying page ranges, thereby substantiating the efficacy of these components.
Effectiveness of the AutoThink Router.
Our arbitrary query router is implemented using a 3B Small Language Model (SLM). To evaluate its contribution, we conduct an ablation study where the router is removed while the subsequent decoupled reasoning module is retained. We report the distribution of hypergraphs across varying page ranges (Figure 3) and the accuracy across diverse scenarios (Table 2). The results demonstrate that the SLM router significantly reduces the proportion of hypergraphs, an effect that becomes more pronounced as the document length increases. This indicates that the router effectively mitigates information construction complexity during initial query processing. Consequently, it yields accuracy gains by simplifying complex reasoning tasks, such as multi-hop question answering.
Effectiveness of the Decomposition of Perception and Reasoning.
Our decoupling module transforms the direct reasoning of the VLM into a hybrid VLM+LLM indirect reasoning paradigm. To evaluate its efficacy, we conduct an ablation study by replacing this module with a standalone LLM reasoning component while keeping the upstream router constant. We evaluate the accuracy across various page ranges, as illustrated in Figure 2. The results indicate that direct VLM reasoning substantially underperforms compared to our decoupled approach, with the performance gap becoming more pronounced as the document length increases.
5 Conclusion
We present AutoThinkRAG, a multi-model collaborative framework that combines a Query Complexity Router with a functionally-decoupled pipeline to tackle long-context and information-overload challenges in information-intensive DocQA. By leveraging a lightweight VLM as a high-fidelity visual interpreter and an LLM for logical deduction, AutoThinkRAG sets new state-of-the-art results on DocBench and MMLongBench while significantly cutting inference costs. Extensive experiments and ablations confirm the effectiveness of both the router and the decoupled architecture.
Limitations
While Auto-thinkRAG achieves significant performance gains over existing multimodal GraphRAG architectures using models with fewer parameters, it still relies on a sequential document parsing and embedding pipeline, which limits the overall processing speed. Our future work will be dedicated to achieving the efficient coupling of document parsing and information encoding to improve the overall processing speed and accuracy.
Ethical Considerations
All our experiments are conducted using widely adopted and recognized open-source datasets curated for academic research purposes. Consequently, these datasets do not contain any inappropriate or sensitive information. While our framework achieves state-of-the-art performance, it is important to acknowledge that our model also builds upon community-driven open-source models and is not flawless. Therefore, users should exercise caution and not rely exclusively on the generated results. In real-world applications, human intervention is recommended to ensure accuracy.

### 17 Mar 2026 v2

@@
 Abstract
-Information-intensive Document Question Answering (DocQA) is often constrained by long contexts and information overload, which hinders Vision-Language Models (VLMs) from performing precise direct reasoning.
-Although multimodal GraphRAG has achieved preliminary breakthroughs, existing frameworks still face dual challenges: (1) the necessity of large-scale models for handling queries of diverse complexities and (2) the inherent reasoning bottlenecks of end-to-end VLMs.
-To address these issues, we propose AutoThinkRAG, a framework that enhances the understanding of complex documents by synergizing the capabilities of multiple models.
-Specifically, we introduce a Query Complexity Router to allocate reasoning paths based on the analysis of query difficulty.
-Furthermore, to overcome the reasoning boundaries of VLM, we propose a functional decoupling architecture: a small-scale VLM serves as a high-fidelity visual interpreter to transform query-relevant visual cues into textual representations, which are subsequently processed by an LLM for logical deduction and synthesis.
-Extensive experiments on DocBench and MMLongBench demonstrate that AutoThinkRAG significantly reduces inference costs while achieving new state-of-the-art performance.
-Further ablation studies verifies the effectiveness of our proposed method.
+Multimodal document question answering requires retrieving dispersed evidence from visually rich long documents and performing reliable reasoning over heterogeneous information.
+Existing multimodal RAG systems remain limited by two bottlenecks: static retrieval that ignores query complexity, and end-to-end Vision-Language Models (VLMs) that couple visual perception with logical reasoning, leading to inefficient computation and unstable answer generation.
+We propose AutoThinkRAG, a complexity-aware inference architecture for multimodal document QA.
+It has two components: (1) a Query Complexity Router that analyzes query difficulty and structure to adaptively select retrieval and reasoning paths; and (2) a Perception–Reasoning Decoupling architecture that uses a lightweight VLM as a high-fidelity visual interpreter to convert query-relevant visual cues into textual representations, which are then passed to an LLM for logical reasoning and answer synthesis.
+This design improves both efficiency and robustness, especially on long-document and unanswerable queries.
+Experiments on DocBench and MMLongBench show that AutoThinkRAG achieves 82.13% and 51.29% overall accuracy, respectively, while reducing per-query token consumption by 18.9% and monetary cost by 18.2%.
+Further analyses show that the gains are most pronounced on complex queries requiring adaptive retrieval and multi-step reasoning.
 AutothinkRAG: Complexity-Aware Control of Retrieval-Augmented Reasoning for Image-Text Interaction
@@
 1 Introduction
-Retrieval-Augmented Generation (RAG) serves as a core technology for processing heterogeneous multimodal documents, including financial PDFs and technical diagrams.
+Figure 1: Compared with traditional RAG, AutothinkRAG enables more effective and robust long-document question answering through adaptive retrieval and decoupled perception-and-reasoning.
+Retrieval-Augmented Generation (RAG) serves as a core technology for processing heterogeneous multimodal documents.
 Its objective is to retrieve multimodal information relevant to a given query from documents as contextual content for question answering, thereby enhancing the accuracy and richness of generated results.
 As an evolving direction in this field, Multimodal RAG (MM-RAG) needs to address two core issues simultaneously: multimodal information processing and high-precision reasoning.
-Knowledge graphs can effectively represent the relationships and attributes of information, while leveraging dense vector representations to identify semantically relevant content lacking direct structural connections.
+Knowledge graphs can effectively represent the relationships and attributes of information, while dense vector representations can identify semantically relevant content lacking direct structural connections.
 In recent years, researchers have conducted extensive explorations on multimodal graph RAG and complex question answering.
-As the state-of-the-art (SOTA) approach, RAGAnything (Guo et al., 2025) has validated the application of multimodal graphs in document question answering.
+As the state-of-the-art (SOTA) approach, RAG-Anything (Guo et al., 2025) has validated the application of multimodal graphs in document question answering.
 HyperGraphRAG (Luo et al., 2025) effectively captures high-order dependencies in complex knowledge bases through hyperedge modeling, improving the accuracy of complex multi-hop reasoning.
-MMGraphRAG (Wan and Yu, 2025) provides a new technical paradigm for multimodal RAG in graph RAG scenarios.
 Despite the progress achieved by existing methods, current SOTA frameworks still face two critical issues that severely impact performance: (1) Retrieval Rigidity: Existing systems adopt a static retrieval strategy that fails to make reasonable judgments and planning based on query complexity.
-Accurate understanding of queries with arbitrary complexity requires large-scale models, and this high-computation-cost approach leads to the inefficient allocation of computing resources.(2) Reasoning Deficit: Multimodal information question answering relies on monolithic Vision-Language Models (VLMs) for end-to-end generation.
-However, recent related studies have shown that VLMs exhibit significantly lower reasoning performance than Large Language Models (LLMs), which consequently results in the problem of "correct visual recognition but incorrect answer generation"(Wan and Yu, 2025).
-In this paper, we propose the AutoThink-RAG framework, which addresses the aforementioned issues through modular cognitive orchestration.
+Accurate understanding of queries with arbitrary complexity requires large-scale models, and this high-computation-cost approach leads to inefficient allocation of computing resources.
+(2) Reasoning Deficit: Multimodal information question answering relies on monolithic Vision-Language Models (VLMs) for end-to-end generation.
+However, recent studies have shown that VLMs exhibit significantly lower reasoning performance than Large Language Models (LLMs), leading to the problem of “correct visual recognition but incorrect answer generation” (Wan and Yu, 2025).
+In this paper, we propose the AutothinkRAG framework to address the aforementioned issues.
 To tackle retrieval rigidity, we design the Autothink lightweight cognitive router, which dynamically selects retrieval paths based on query parsing and complexity evaluation, enabling on-demand allocation of computing resources.
-To resolve reasoning deficit, we propose the Decomposition of Perception and Reasoning (DPR) architecture, which limits the role of VLMs to that of "visual translators" and delegates logical integration tasks to LLMs, achieving more accurate reasoning without loss of relevant information.
-Our contributions are three-fold:
-  • We present AutoThink-RAG, a scalable architecture that integrates MinerU-based parsing with a hybrid Graph-Vector storage, establishing a new pareto-optimal frontier between efficiency and accuracy;
+To resolve reasoning deficit, we propose the Decomposition of Perception and Reasoning (DPR) architecture, which limits the role of VLMs to that of “visual descriptors” and delegates logical integration tasks to LLMs, achieving more accurate reasoning without loss of relevant information.
+Our contributions are four-fold:
+  • We present AutothinkRAG, a scalable architecture that integrates MinerU-based parsing with a hybrid Graph-Vector storage, establishing a new pareto-optimal frontier between efficiency and accuracy;
   • We design the Autothink Router to handle queries of unknown complexity; by leveraging a lightweight Small Language Model (SLM) for in-depth complexity analysis and task decomposition, this module enables adaptive execution path selection, effectively addressing the challenge of Retrieval Rigidity.
@@
 Beyond binary relations, hypergraph and higher-arity modeling improves expressiveness and reduces information loss in complex reasoning (HyperGraphRAG (Luo et al., 2025), Structure Is All You Need (Lee and Whang, )).
+Figure 2: Overview of our AutothinkRAG.
+The framework consists of four primary stages: (1) Information Extraction: Heterogeneous multimodal data are ingested.
+(2) Data Processing: Through layout analysis, the system partitions data into text chunks and image indices, followed by an iterative “Extract–Judge” loop to construct a Graph Knowledge Base.
+(3) Data Linker: A hybrid retrieval mechanism integrates traditional Top–K vector search with graph–based merging to capture complex entity relationships.
+(4) Reasoning & Response: The Query Router decomposes the original query into sub-queries ($Q_{1},Q_{2},Q_{3}$), which are then processed through a decoupled perception reasoning architecture leveraging both “related text” and visual cues to synthesize the final response.
 2.2 Multimodal Retrieval-Augmented Generation
@@
 3.1 Overall Workflow
-AutoThinkRAG deconstructs the multimodal Document Question Answering (DocQA) task into intent-aware routing and decoupled reasoning. The end-to-end inference process is formally defined as:
+AutothinkRAG deconstructs the multimodal Document Question Answering (DocQA) task into intent-aware routing and decoupled reasoning,as illustrated in Figure  2.
+The end-to-end inference process is formally defined as:
 $A=\Psi\left(Q,\mathcal{D}\right)=\text{LLM}\left(T_{v},\mathcal{R},I_{p}\right)$ (1)
@@
   2. 2.
-    Vector Store: All entities, relations, and text chunks are mapped into a dense embedding space, with $S,B,$ and $P$ indexed as key attributes.
+    Vector Store: All entities, relations, and text chunks are mapped into a dense embedding space, with $S,B,$ and $P$ indexed as attributes.
 Information Transmission Mechanism.
-The framework achieves efficient cross-module transmission via a metadata-driven protocol.
-During retrieval, the storage path $S$ enables the direct transmission of raw multimodal data to the VLM.
-Simultaneously, leveraging $B$ and $P$, the system fetches adjacent context from the original document, effectively bridging the “Information Gap” between isolated fragments and their original document context.
-Doc Types Domains All
-Aca. Fin. Gov. Law. News Txt. Mm. Una.
-LLM 41.3 16.3 50.7 49.7 77.3 53.9 20.1 75.8 44.8
-MLLM 19.8 16.3 28.4 31.4 46.5 35.7 15.9 39.5 27.7
-$M$ Locations Formats Overall
-Sin. Mul. Una. C.T. Txt. Lay. Fig. All
-LLM 22.5 20.0 53.2 14.7 33.3 23.5 16.1 27.8
-MLLM 13.3 7.9 13.9 8.4 10.7 11.8 11.4 11.6
-Table 1: Comparative experimental results of VLM and LLM on DocBench and MMLongBench benchmarks.
-Figure 1: Overview of our AutoThink–RAG.
-The framework consists of four primary stages: (1) Information Extraction: Heterogeneous multimodal data are ingested.
-(2) Data Processing: Through layout analysis, the system partitions data into text chunks and image indices, followed by an iterative “Extract–Judge” loop to construct a Graph Knowledge Base.
-(3) Data Linker: A hybrid retrieval mechanism integrates traditional Top–K vector search with graph–based merging to capture complex entity relationships.
-(4) Reasoning & Response: The Query Router decomposes the original query into sub-queries ($Q_{1},Q_{2},Q_{3}$), which are then processed through a decoupled perception reasoning architecture leveraging both “related text” and visual cues to synthesize the final response.
+The framework enables efficient cross-module transmission via a metadata-driven protocol.
+During retrieval, storage path $S$ allows direct raw multimodal data transmission to the VLM.
+Meanwhile, leveraging $B$ and $P$, the system fetches adjacent context from the original document, effectively bridging the “Information Gap” between isolated fragments and their source context.
 3.3 Autothink Router and Functional Decoupling Architecture
@@
     Dependency Feature: Identifies cross-chunk/multi-step reasoning needs ($\text{ctx}(Q)$).
-Based on these, QCR assigns a complexity label $c(Q)$ and generates routing instructions $I_{p}$.
+Based on these, QCR assigns a complexity label $c(Q)$ and generates routing instructions $I_{p}$.Concretely, queries relying on localized evidence are regarded as Simple, whereas queries requiring multi-entity, multimodal aggregation and multi-step reasoning are regarded as Complex.
+Doc Types Domains All
+Aca. Fin. Gov. Law. News Txt. Mm. Una.
+LLM 41.3 16.3 50.7 49.7 77.3 53.9 20.1 75.8 44.8
+MLLM 19.8 16.3 28.4 31.4 46.5 35.7 15.9 39.5 27.7
+$M$ Locations Formats Overall
+Sin. Mul. Una. C.T. Txt. Lay. Fig. All
+LLM 22.5 20.0 53.2 14.7 33.3 23.5 16.1 27.8
+MLLM 13.3 7.9 13.9 8.4 10.7 11.8 11.4 11.6
+Table 1: Comparative experimental results of VLM and LLM on DocBench and MMLongBench benchmarks.
+Method Domains Types Overall
+Aca. Fin. Gov. Law. News. Txt. Mm. Una.
+AutothinkRAG 80.16 75.40 85.71 87.09 92.50 86.43 77.31 81.25 82.13
+Nothink 77.34 73.46 80.59 85.29 88.78 84.08 76.12 77.84 79.37
+VLM only 77.03 74.00 81.54 84.39 89.26 85.32 75.09 76.42 79.47
+baseline 76.34 72.93 79.77 84.07 84.52 84.65 74.47 75.83 78.02
+RAG-Anything 74.90 70.68 77.41 83.06 81.67 84.21 76.36 52.80 75.47
+Table 2: Accuracy (%) on DocBench Dataset.
+Domain categories include Academia (Aca.), Finance (Fin.), Government (Gov.), Legal Documents (Law), and News Articles (News).
+Document types are categorized as Text-only (Txt.), Multimodal (Mm.), and Unanswerable queries (Una.).
+Method Domains Overall
+Res. Tut. Acad. Guid. Broch. Admin. Fin.
+AutothinkRAG 44.86 52.20 52.26 52.34 51.00 62.19 56.14 51.29
+Nothink 43.15 48.55 43.21 52.25 46.00 53.08 50.42 47.21
+VLM (direct) 41.43 49.27 43.43 52.25 41.00 54.32 48.72 45.55
+VLM (unguided) 42.07 51.12 45.24 52.34 43.00 55.78 50.13 47.21
+baseline 40.41 47.82 42.42 50.96 42.00 51.85 46.15 44.86
+8B 42.80 44.12 43.21 52.34 46.00 50.61 50.42 46.09
+RAG-Anything 41.09 41.30 44.22 45.16 40.00 54.32 52.13 44.36
+Table 3: Accuracy (%) on MMLongBench across different domains and overall performance.
+Domain categories include Research Reports/Introductions (Res.), Tutorials/Workshops (Tut.), Academic Papers (Acad.), Guidebooks (Guid.), Brochures (Broch.), Administration/Industry Files (Admin.), and Financial Reports (Fin.).
 3.3.2 Functional Decoupling via DPR
-Recent studies indicate that VLMs exhibit significantly lower capabilities in logical reasoning compared to LLMs (as shown in Table 1).
-This discrepancy stems from the VLM architecture’s over-reliance on intermediate projection layers for modality alignment, which struggles to maintain logical consistency in complex derivations.
-Conversely, VLMs demonstrate superior performance in descriptive tasks post-perception.
+Recent studies indicate that VLMs exhibit significantly lower capabilities in logical reasoning compared to LLMs, particularly in document question answering tasks involving unanswerable queries.
+As reported in MMGraphRAG Wan and Yu (2025), LLM-based systems consistently outperform multimodal models on the DocBench and MMLongBench benchmarks.
+Notably, VLM-based models tend to produce hallucinated answers when the required evidence is absent (as shown in Table 1).
+In long-document multimodal QA, visual perception, cross-page evidence alignment, and reasoning are tightly coupled within a single generation, making the model prone to cascading errors: weakly aligned evidence causes overcommitment to unsupported answers instead of reliable abstention.Conversely, VLMs demonstrate superior performance in descriptive tasks post-perception.
 Consequently, we propose the Decoupled Perception and Reasoning (DPR) strategy, partitioning the task into two specialized sub-processes:
@@
 By decoupling perception from reasoning, the LLM performs rigorous deduction over precise textualized visual evidence, effectively bridging the reasoning gap of standalone multimodal models.
-Method Domains Types Overall
-Aca. Fin. Gov. Law. News. Txt. Mm. Una.
-AutothinkRAG 80.16 75.40 85.71 87.09 92.50 86.43 77.31 81.25 82.13
-Nothink 77.34 73.46 80.59 85.29 88.78 84.08 76.12 77.84 79.37
-VLM only 77.03 74.00 81.54 84.39 89.26 85.32 75.09 76.42 79.47
-baseline 76.34 72.93 79.77 84.07 84.52 84.65 74.47 75.83 78.02
-Raganything 74.90 70.68 77.41 83.06 81.67 84.21 76.36 52.80 75.47
-Table 2: Accuracy (%) on DocBench Dataset.
-Domain categories include Academia (Aca.), Finance (Fin.), Government (Gov.), Legal Documents (Law), and News Articles (News).
-Document types are categorized as Text-only (Txt.), Multimodal (Mm.), and Unanswerable queries (Una.).
-Method Domains Overall
-Res. Tut. Acad. Guid. Broch. Admin. Fin.
-AutothinkRAG 44.86 52.20 52.26 52.34 51.00 62.19 56.14 51.29
-Nothink 43.15 48.55 43.21 52.25 46.00 53.08 50.42 47.21
-VLM only 41.43 49.27 43.43 52.25 41.00 54.32 48.72 45.55
-baseline 40.41 47.82 42.42 50.96 42.00 51.85 46.15 44.86
-8B 42.80 44.12 43.21 52.34 46.00 50.61 50.42 46.09
-Raganything 41.09 41.30 44.22 45.16 40.00 54.32 52.13 44.36
-Table 3: Accuracy (%) on MMLongBench across different domains and overall performance.
-Domain categories include Research Reports/Introductions (Res.), Tutorials/Workshops (Tut.), Academic Papers (Acad.), Guidebooks (Guid.), Brochures (Broch.), Administration/Industry Files (Admin.), and Financial Reports (Fin.).
 4 Experiments
-In this section, we empirically validate the effectiveness of AutoThink-RAG.
+Figure 3: Comparison of routing strategy distributions across different document lengths.
+This figure illustrates the selection proportions of graphs and hypergraphs within length intervals under Think and No-Think modes.
+In this section, we empirically validate the effectiveness of AutothinkRAG.
 Our goal is to address two primary research questions: (1) Compared to static retrieval, can the Autothink router effectively reduce hallucinations and handle complex query intents?
@@
 4.2 Implementation Details
-Figure 2: Comparison of routing strategy distributions across different document lengths.
-This figure illustrates the selection proportions of graphs and hypergraphs within length intervals under Think and No-Think modes.
-In our experimental evaluation of AutoThink-RAG, we implement a multi-model collaborative framework characterized by a decoupled architecture.
+In our experimental evaluation of AutothinkRAG, we implement a multi-model collaborative framework characterized by a decoupled architecture.
 Specifically, we employ Qwen2.5-VL-3B as the Autothink Router to facilitate global task orchestration and intent scheduling.
@@
 Performance on DocBench.
-Table 2 summarizes the performance on DocBench. AutoThink-RAG achieves a state-of-the-art Overall Accuracy of 82.13%, significantly outperforming the baseline (78.02%).
-Crucially, our framework demonstrates a substantial breakthrough in the Unanswerable (Una) category, boosting accuracy from 52.80% to 81.25% (a +28.45% improvement).
-This result strongly validates the cognitive awareness of our AutoThink router.
-Unlike static retrieval systems that force-feed irrelevant context to the generator, AutoThink can detect insufficient information and guide the reasoning engine to refuse answering, thereby drastically reducing hallucinations.
-Additionally, in information-dense domains like News (+10.83%) and Government (+8.30%), our Hybrid Hypergraph Retrieval proves effective in capturing complex entity relationships that simple vector search misses.
+Table 2 shows that AutothinkRAG achieves the best Overall Accuracy of 82.13%, outperforming the baseline (78.02%).
+The largest gain appears in Unanswerable (Una), where accuracy improves from 52.80% to 81.25% (+28.45%), indicating the effectiveness of the Autothink router in detecting insufficient evidence and reducing hallucinations.
+Consistent improvements on News (+10.83%) and Government (+8.30%) further demonstrate the advantage of Hybrid Hypergraph Retrieval in modeling complex entity relations beyond simple vector search.
 Figure 4: The approach of converting VLM outputs to LLM leads to a marked boost in accuracy when processing multi-page long documents.
 Performance on MMLongBench.
-Table 3 presents the results on long-context tasks. AutoThink-RAG consistently surpasses the baseline across all document categories, achieving an Overall Accuracy of 51.29% (+6.43%).
-The unified baseline often struggles with "contextual entrainment" in long documents, where visual noise distracts the VLM’s reasoning.
-In contrast, our DPR Architecture maintains high performance even in complex categories like Admin (+10.34%) and Finance (+9.99%).
-By converting visual signals into dense text descriptions and offloading the synthesis to a specialized LLM, our approach effectively bridges the "Reasoning Gap," allowing for robust multi-hop deduction over extensive contexts.
+Table 3 reports the results on long-context tasks.
+AutothinkRAG consistently outperforms the baseline across all document categories, achieving an Overall Accuracy of 51.29% (+6.43%).
+While the baseline struggles with visual noise in long documents, our DPR Architecture remains effective in complex categories such as Admin (+10.34%) and Finance (+9.99%), bridging the “Reasoning Gap” by converting visual signals into textual descriptions and delegating synthesis to a specialized LLM.
+To quantify the efficiency of AutothinkRAG, we analyze per-query token consumption and monetary cost (Table 4 and Table 5).
+The overall inference comprises three stages: query routing (Qwen2.5-VL-3B), visual description generation (Qwen2.5-VL-7B), and final answer synthesis (LLM).
+The input context of the synthesis stage accounts for the dominant cost.
+Without routing, pure hypergraph retrieval feeds substantial redundant context into the LLM, costing 63.0k tokens / 0.132 RMB per query.
+With the Autothink router, the synthesis input is significantly reduced to 51.4k tokens / 0.108 RMB, yielding savings of 18.9% and 18.2%, respectively.
+This demonstrates that the router effectively avoids unnecessary long-context reasoning by dynamically selecting retrieval paths based on query complexity.
 4.4 Baselines
-To ensure experimental rigor and fairness, we built a baseline model and employed the identical model on RAGAnything (Guo et al., 2025) for testing and comparison.In the baseline configuration, we bypass query decomposition and employ the VLM for direct end-to-end inference, rather than using our proposed decoupled paradigm.
-For the RagAnything benchmark, we substitute its internal components with the specific models used in our framework and maintain identical hyperparameter configurations to ensure a controlled comparison.
+To ensure experimental rigor and fairness, we built a baseline model and employed the identical model on RAG-Anything (Guo et al., 2025) for testing and comparison.In the baseline configuration, we bypass query decomposition and employ the VLM for direct end-to-end inference, rather than using our proposed decoupled paradigm.
+For the RAG-Anything benchmark, we substitute its internal components with the specific models used in our framework and maintain identical hyperparameter configurations to ensure a controlled comparison.
 4.5 Ablation Study
@@
 Consequently, it yields accuracy gains by simplifying complex reasoning tasks, such as multi-hop question answering.
+Pipeline Image Description Summary Total
+Hypergraph 2.7k 60.3k 63.0k
+Routing-Enhanced 2.7k 48.3k 51.4k
+Savings (Abs.) - 12.0k 11.6k
+Savings (Rel.) - 19.9% 18.9%
+Table 4: Token comparison across stages.
+Pipeline Image Description Summary Total
+Hypergraph 0.0084 0.1239 0.132
+Routing-Enhanced 0.0084 0.0984 0.108
+Savings (Abs.) - 0.0255 0.024
+Savings (Rel.) - 20.6% 18.2%
+Table 5: Monetary cost comparison across stages.
 Effectiveness of the Decomposition of Perception and Reasoning.
@@
 To evaluate its efficacy, we conduct an ablation study by replacing this module with a standalone LLM reasoning component while keeping the upstream router constant.
-We evaluate the accuracy across various page ranges, as illustrated in Figure 2.
-The results indicate that direct VLM reasoning substantially underperforms compared to our decoupled approach, with the performance gap becoming more pronounced as the document length increases.
+We evaluate the accuracy across various page ranges, as illustrated in Figure  4.
+The results indicate that both direct VLM reasoning and unguided VLM decoupling for visual perception in long-document settings substantially underperform the proposed full decoupling framework, with the performance gap becoming more pronounced as the document length increases.
 5 Conclusion
@@
 Limitations
-While Auto-thinkRAG achieves significant performance gains over existing multimodal GraphRAG architectures using models with fewer parameters, it still relies on a sequential document parsing and embedding pipeline, which limits the overall processing speed.
+While AutothinkRAG achieves significant performance gains over existing multimodal GraphRAG architectures using models with fewer parameters, it still relies on a sequential document parsing and embedding pipeline, which limits the overall processing speed.
 Our future work will be dedicated to achieving the efficient coupling of document parsing and information encoding to improve the overall processing speed and accuracy.